# Quantum Spin Liquids in Pyrochlore Magnets With Non-Kramers Local Moments

### v1

@@ I Introduction
Quantum spin liquids (QSLs) are novel quantum paramagnetic states of matter produced by long-range entanglement of interacting spins [1, 2, 3, 4, 5]. QSLs do not order at zero temperature, and they possess fractionalized excitations such as spinons and emergent gauge fields [6, 7, 8, 9, 10, 11, 12, 13]. While a dimer liquid form of QSL is realized in Rydberg atom arrays [14, 15], the discovery of QSL in real materials remains an outstanding challenge in quantum materials research.
Pyrochlore magnets have been considered as possible platforms for QSLs due to the geometric frustration of the pyrochlore lattice (Fig. 1 (a)) [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. In particular, several promising signatures of QSL behaviour have been observed in pyrochlore rare-earth oxides $\textrm{Pr}_{2}(\textrm{Zr, Sn, Hf, Ir})_{2}\textrm{O}_{7}$ [31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. These compounds have the interesting property that the magnetically active $\textrm{Pr}^{3+}$ ions are described as a non-Kramers doublet at low energies, which is protected by crystal symmetries, but not by Kramers degeneracy [30]. The non-Kramers doublet is conveniently represented by pseudospins-$1/2$, where the $S^{z}$ component transforms as a dipole but the local $S^{x}$ and $S^{y}$ components transform as quadrupoles [30, 41]. Heat capacity measurements down to $0.35$ K and DC magnetic susceptibility measurements down to $1.8$ K show no evidence of a phase transition [31, 32, 34, 35, 37, 40]. Furthermore, neutron scattering experiments in $\textrm{Pr}_{2}\textrm{Zr}_{2}\textrm{O}_{7}$, $\textrm{Pr}_{2}\textrm{Hf}_{2}\textrm{O}_{7}$, and $\textrm{Pr}_{2}\textrm{Sn}_{2}\textrm{O}_{7}$ find an inelastic continuum of excitations that could be attributed to fractionalized quasiparticles of a QSL [31, 32, 42, 38, 34, 35].
It has been suggested that the putative QSL in $\textrm{Pr}_{2}\textrm{Zr}_{2}\textrm{O}_{7}$ and $\textrm{Pr}_{2}\textrm{Hf}_{2}\textrm{O}_{7}$ is quantum spin ice (QSI) [31, 43, 38] with dominant dipolar Ising exchange and smaller quadrupolar transverse exchange interactions. The minimal model for QSI is the XXZ model on the pyrochlore lattice
@@
where $J_{zz}>0$. In classical spin ice (i.e., $J_{\pm}=0$), the ground state spin configuration is characterized by the local two-in-two-out constraint, resulting in a macroscopic degeneracy [44]. An excited state with energy gap $J_{zz}$ can be produced by flipping one spin in a particular ground state, creating two “defect tetrahedra” which break the two-in-two-out constraint [44, 45, 46]. Going to the quantum theory in the perturbative regime $J_{\pm}\ll J_{zz}$, the action of the transverse operators $S^{\pm}$ on the ground state is to create a pair of dispersive spinon-antispinon excitations with an excitation gap on the order of $J_{zz}/2$. Formally, QSI can be described by mapping pseudospins to lattice quantum electrodynamics [16, 19, 47, 48, 49]
@@
where $i$ labels the sites of the pyrochlore lattice and $r,r^{\prime}$ represent the centers of the tetrahedra sharing the site $i$. Here, $\mathcal{E}_{rr^{\prime}}$ is a lattice electric field and $\Phi^{\dagger}_{r}\Phi_{r^{\prime}}$ represents the creation of a spinon and antispinon pair. Physically, the spinons (“electric charges”) are the quantum analog of the classical defect tetrahedra, while fluctuations of the electric field describe gapless photon excitations of the QSI [16, 19, 47, 48, 49, 50].
@@
In the case of Pr-based pyrochlore magnets, if the ground state is a QSI with the dominant dipolar Ising exchange interaction, only the dipolar Ising moment $S^{z}$ or the emergent electric field couples linearly to the neutron spin. On the other hand, the gapped excitations of QSI involve $S^{\pm}$ quadrupolar moments and hence they do not couple to neutron spins. Furthermore, although the dual magnetic monopoles or “vison” excitations can couple to neutrons in principle [51], detailed calculations show that such a signal is several orders of magnitude weaker than any contribution from the photons [52]. Hence, neutron scattering is expected to detect the quasielastic low-energy photons while being essentially insensitive to the inelastic signals from gapped spinon excitations and the extremely weak magnetic monopole contribution. On the other hand, neutron scattering in $\textrm{Pr}_{2}\textrm{Zr}_{2}\textrm{O}_{7}$ and $\textrm{Pr}_{2}\textrm{Hf}_{2}\textrm{O}_{7}$ shows significant spectral weight in the inelastic signals, which is inconsistent with the interpretation of the QSI [31, 32, 42, 53, 38]. Interestingly, the inelastic signals show “rod”-like structures in the $[hhl]$ scattering plane [54]. It has been suggested that these inelastic signals may be due to disorder effects [55, 33, 42, 56, 57]. Finally, there is no consensus on the exchange parameters of $\textrm{Pr}_{2}(\textrm{Zr, Sn, Hf, Ir})_{2}\textrm{O}_{7}$, so we have no clear reason to assume that they lie in or even near the spin ice regime in parameter space even without disorder.
@@
Given these facts, we explore alternative quantum spin liquid candidates in pyrochlore magnets with non-Kramers local moments. In this work, instead of considering disorder effects, we investigate the feasibility of quantum spin liquids beyond the quantum spin ice regime, which may explain the dominant inelastic signals in neutron scattering experiments. We study possible $\mathbb{Z}_{2}$ QSLs using Schwinger boson mean-field theory, restricting our attention to QSLs which do not break any space group symmetries or time-reversal symmetry. We use the projective symmetry group (PSG) to classify all such QSLs [58, 59, 60]. We construct the phase diagram and identify the spin liquid candidates that are consistent with the experiments. Among them, the most robust candidates occur in the region where the transverse quadrupolar exchange coupling $J_{\pm}$ is dominant. This is a parameter regime where studies on the XXZ model (1) have identified a $U(1)$ nematic QSL through exact diagonalization and variational calculations [61] and a disordered state with a large nematic susceptibility in pseudo-Majorana fermion functional renormalization group [62]. The $\mathbb{Z}_{2}$ QSLs identified here might be closely competing with these disordered states and may be stabilized by couplings beyond the pristine XXZ model. Finally, we compute the static and dynamic spin structure factors for these candidates to identify their experimental signatures in neutron scattering.
@@
Figure 1: (a) The pyrochlore lattice and global Cartesian coordinate frame, where “up-pointing” and “down-pointing” tetrahedra are coloured green and blue, respectively. (b) An up-pointing tetrahedron with local $\hat{z}$ axes and sublattice indices attached. (b) Phase diagram in the region $J_{\pm}<0$ with $J_{zz}^{2}+J_{\pm}^{2}+J_{\pm\pm}^{2}=1$ indicating the lowest energy PSG classes. We set $\kappa=0.1$. The XXZ model occurs along the horizontal line joining $\phi=0$ to $\phi=\pi$. The point H indicates the antiferromagnetic Heisenberg point $J_{\pm}/J_{zz}=-\frac{1}{2},J_{\pm\pm}=0$. The $\times$’s indicate the points at which the structure factors were calculated for the relevant PSG classes (see table 1). (c) Phase diagram at $\kappa=0.1$ indicating the second-lowest energy PSG classes.
@@ II Model
[…]
@@
[…]
where $\alpha,\beta,\gamma\in\mathbb{R}$ [63]. The resulting $S^{z}$ component of the non-Kramers doublet transforms as a dipole, while the in-plane $S^{x},S^{y}$ components transform as quadrupoles [21].
@@
  • $S$: a “screw” operation defined by fractional translation by $\hat{\epsilon}_{3}=\frac{1}{2}{\bm{\hat{e}}}_{3}$ composed with a $\pi$ rotation around ${\bm{\hat{e}}}_{3}$. [41, 60, 64]
Finally, it is convenient to work in sublattice-dependent local bases to express the orientations of the pseudospins. We attach an orthonormal basis $\{\hat{x}_{\mu},\hat{y}_{\mu},\hat{z}_{\mu}\}$ to each sublattice $\mu$ (expressions given in table 2). As shown in Fig. 1 (b), these have the property that each $\hat{z}_{\mu}$ basis vector points out of the “up-pointing” tetrahedra.
@@
where $i,j$ denote pyrochlore lattice sites, and $\gamma_{ij}$ is a bond-dependent phase factor defined in appendix A [21]. We take $J_{\pm\pm}>0$ without loss of generality because its sign can be flipped by a $\frac{\pi}{2}$ rotation of each local basis $\{\hat{x}_{\mu},\hat{y}_{\mu},\hat{z}_{\mu}\}$ about $\hat{z}_{\mu}$ [65].
@@ III.1 Schwinger Bosons
$S^{a}_{i}$ $=\frac{1}{2}\sum_{\alpha\beta}b^{\dagger}_{i\alpha}[\sigma^{a}]_{\alpha\beta}b_{j\beta}$ (9)
where $a\in\{x,y,z\}$ and $\sigma^{a}$ are the Pauli matrices. The operators $b_{i\alpha}$ satisfy the bosonic commutation relations $[b_{i\alpha},b^{\dagger}_{j\beta}]=\delta_{ij}\delta_{\alpha\beta}$. The Schwinger bosons physically represent the fractionalized excitations of the QSL, which we refer to as spinons. If we impose the local constraint
@@
with $\kappa=1$, we obtain a faithful representation of the original pseudospin-$1/2$ Hilbert space [28, 66]. In the mean-field theory, this constraint is relaxed to only be satisfied on average
@@
for all sites $i$. We will also treat $\kappa$ as a continuous positive parameter, which gives us control over the size of quantum effects. Taking $\kappa\to\infty$ recovers the classical limit, and $\kappa<1$ corresponds to larger quantum fluctuations. In the current work, we focus on small $\kappa\ll 1$ to obtain all the competing QSLs, and comment on what is expected in the physical $\kappa=1$ limit later in sections IV.2 and V [67]. As $\kappa$ increases, magnetic order may develop in some parts of the phase diagram, signaled by the closing of the spinon gap and condensation of Schwinger bosons [66, 68, 60].
@@ III.2 Mean-Field Decoupling
[…]
@@ III.3 Projective Symmetry Group Classification
Two ansätze which differ by a gauge transformation of the form (19) are therefore physically equivalent [58, 8, 59]. The correct way to implement symmetry operations is via the gauge-enriched operations $\tilde{{\mathcal{O}}}=G_{\mathcal{O}}\circ{\mathcal{O}}$, where ${\mathcal{O}}$ is an element of the pyrochlore lattice space group [58, 59]. Explicitly, the action of $\tilde{{\mathcal{O}}}$ is
@@
with $\mathcal{K}$ being the complex conjugation operator and $U_{\mathcal{T}}$ is a $2\times 2$ unitary matrix [60]. Explicit expressions for $U_{\mathcal{O}}$ and $U_{\mathcal{T}}$ are derived in appendix B; here we simply list the results
@@
where $\omega=e^{2\pi i/3}$. The group generated by the gauge-enriched operations $\tilde{{\mathcal{O}}}$ and $\tilde{\mathcal{T}}$ is known as the projective symmetry group (PSG) [58, 59, 60]. Demanding that the mean-field Hamiltonian is invariant under the PSG elements puts constraints on the allowed mean-field ansätze. Ansätze that cannot be transformed into each other via a pure gauge transformation (19) belong to different PSG classes, meaning they describe different QSLs [58, 59].
The subgroup of the PSG corresponding to pure gauge transformations associated with the identity operation ${\mathcal{O}}=\mathrm{id}$ is known as the invariant gauge group [58, 59]. The invariant gauge group corresponds to the emergent gauge structure of the resulting QSLs [58, 8]. In this paper, we only consider the $\mathbb{Z}_{2}=\{e^{i\pi n}|n\in\{0,1\}\}$ invariant gauge group, thus restricting our study to $\mathbb{Z}_{2}$ symmetric bosonic QSLs.
@@
where $n_{1},n_{\overline{C}_{6}S},n_{ST_{1}},n_{\overline{C}_{6}}$ are all either 00 or $1$ [60]. We do not consider the case $n_{1}=1$, which has an enlarged unit cell because the translations $T_{2}$ and $T_{3}$ act projectively. We therefore label the different PSG classes using the notation $(n_{\overline{C}_{6}S}n_{ST_{1}}n_{\overline{C}_{6}})$ in this work. Even though the PSG solution is the same as the effective spin-$1/2$ [60] and dipolar-octupolar cases [64], the structure of the allowed mean-field ansätze will be different because of the pseudospin transformation matrices $U_{\mathcal{O}}$ and $U_{\mathcal{T}}$. Following [60], we shift the first Brillouin zone ${{\bm{q}}}\mapsto{{\bm{q}}}-(\pi,\pi,\pi)$ for the PSG classes with ${n_{ST_{1}}}=1$, which is equivalent to a gauge transformation and does not affect physical quantities on the spin level.

### v2

@@ I Introduction
Quantum spin liquids (QSLs) are novel quantum paramagnetic states of matter produced by long-range entanglement of interacting spins [Savary2016, wen2017colloquium, knolle2019field, zhou2017quantum, Broholm2020]. QSLs do not order at zero temperature, and they possess fractionalized excitations such as spinons and emergent gauge fields [Ioffe1989, wen1990topological, wen2004quantum, levin2005colloquium, chen2010local, wen2013topological, wen2017colloquium, wen2019choreographed]. While a dimer liquid form of QSL is realized in Rydberg atom arrays [Verresen2021, Semeghini2021], the discovery of QSL in real materials remains an outstanding challenge in quantum materials research.
Pyrochlore magnets have been considered as possible platforms for QSLs due to the geometric frustration of the pyrochlore lattice (Fig. 1 (a)) [hermele2004pyrochlore, huse2003coulomb, jaubert2011analysis, banerjee2008unusual, shannon2012quantum, Onoda2011, huang2018dynamics, huang2020extended, henley2010coulomb, bramwell2020history, castelnovo2012spin, Gardner2010, Book2011, udagawa2021spin, Rau2019]. In particular, several promising signatures of QSL behaviour have been observed in pyrochlore rare-earth oxides $\textrm{Pr}_{2}(\textrm{Zr, Sn, Hf, Ir})_{2}\textrm{O}_{7}$ [Kimura2013, Petit2016, Martin2017, Zhou2008, Ortiz2024, hatnean2016single, Sibille2016, Sibille2018, Machida2005, Nakatsuji2006]. These compounds have the interesting property that the magnetically active $\textrm{Pr}^{3+}$ ions are described as a non-Kramers doublet at low energies, which is protected by crystal symmetries, but not by Kramers degeneracy [Rau2019]. The non-Kramers doublet is conveniently represented by pseudospins-$1/2$, where the $S^{z}$ component transforms as a dipole but the local $S^{x}$ and $S^{y}$ components transform as quadrupoles [Rau2019, Curnoe2018]. Heat capacity measurements down to $0.35$ K and DC magnetic susceptibility measurements down to $1.8$ K show no evidence of a phase transition [Kimura2013, Petit2016, Zhou2008, Ortiz2024, Sibille2016, Nakatsuji2006]. Furthermore, neutron scattering experiments in $\textrm{Pr}_{2}\textrm{Zr}_{2}\textrm{O}_{7}$, $\textrm{Pr}_{2}\textrm{Hf}_{2}\textrm{O}_{7}$, and $\textrm{Pr}_{2}\textrm{Sn}_{2}\textrm{O}_{7}$ find an inelastic continuum of excitations that could be attributed to fractionalized quasiparticles of a QSL [Kimura2013, Petit2016, Wen2017, Sibille2018, Zhou2008, Ortiz2024].
It has been suggested that the putative QSL in $\textrm{Pr}_{2}\textrm{Zr}_{2}\textrm{O}_{7}$ and $\textrm{Pr}_{2}\textrm{Hf}_{2}\textrm{O}_{7}$ is quantum spin ice (QSI) [Kimura2013, Tokiwa2018, Sibille2018] with dominant dipolar Ising exchange and smaller quadrupolar transverse exchange interactions. The minimal model for QSI is the XXZ model on the pyrochlore lattice
@@
where $J_{zz}>0$. In classical spin ice (i.e., $J_{\pm}=0$), the ground state spin configuration is characterized by the local two-in-two-out constraint, resulting in a macroscopic degeneracy [Gingras2014]. An excited state with energy gap $J_{zz}$ can be produced by flipping one spin in a particular ground state, creating two “defect tetrahedra” which break the two-in-two-out constraint [Gingras2014, bramwell2001spin, castelnovo2008magnetic]. Going to the quantum theory in the perturbative regime $J_{\pm}\ll J_{zz}$, the action of the transverse operators $S^{\pm}$ on the ground state is to create a pair of dispersive spinon-antispinon excitations with an excitation gap on the order of $J_{zz}/2$. Formally, QSI can be described by mapping pseudospins to lattice quantum electrodynamics [hermele2004pyrochlore, banerjee2008unusual, Benton2012, Savary2012, Lee2012]
@@
where $i$ labels the sites of the pyrochlore lattice and $r,r^{\prime}$ represent the centers of the tetrahedra sharing the site $i$. Here, $\mathcal{E}_{rr^{\prime}}$ is a lattice electric field and $\Phi^{\dagger}_{r}\Phi_{r^{\prime}}$ represents the creation of a spinon and antispinon pair. Physically, the spinons (“electric charges”) are the quantum analog of the classical defect tetrahedra, while fluctuations of the electric field describe gapless photon excitations of the QSI [hermele2004pyrochlore, banerjee2008unusual, Benton2012, Savary2012, Lee2012, savary2021quantum].
@@
In the case of Pr-based pyrochlore magnets, if the ground state is a QSI with the dominant dipolar Ising exchange interaction, only the dipolar Ising moment $S^{z}$ or the emergent electric field couples linearly to the neutron spin. On the other hand, the gapped excitations of QSI involve $S^{\pm}$ quadrupolar moments and hence they do not couple to neutron spins. Furthermore, although the dual magnetic monopoles or “vison” excitations can couple to neutrons in principle [chen2017dirac], detailed calculations show that such a signal is several orders of magnitude weaker than any contribution from the photons [kwasigroch2020vison]. Hence, neutron scattering is expected to detect the quasielastic low-energy photons while being essentially insensitive to the inelastic signals from gapped spinon excitations and the extremely weak magnetic monopole contribution. On the other hand, neutron scattering in $\textrm{Pr}_{2}\textrm{Zr}_{2}\textrm{O}_{7}$ and $\textrm{Pr}_{2}\textrm{Hf}_{2}\textrm{O}_{7}$ shows significant spectral weight in the inelastic signals, which is inconsistent with the interpretation of the QSI [Kimura2013, Petit2016, Wen2017, anand2016physical, Sibille2018]. Interestingly, the inelastic signals show “rod”-like structures in the $[hhl]$ scattering plane [Castelnovo2019]. It has been suggested that these inelastic signals may be due to disorder effects [Bonville2016, Martin2017, Wen2017, Savary2017, Benton2018]. Finally, there is no consensus on the exchange parameters of $\textrm{Pr}_{2}(\textrm{Zr, Sn, Hf, Ir})_{2}\textrm{O}_{7}$, so we have no clear reason to assume that they lie in or even near the spin ice regime in parameter space even without disorder.
@@
Given these facts, we explore alternative quantum spin liquid candidates in pyrochlore magnets with non-Kramers local moments. In this work, instead of considering disorder effects, we investigate the feasibility of quantum spin liquids beyond the quantum spin ice regime, which may explain the dominant inelastic signals in neutron scattering experiments. We study possible $\mathbb{Z}_{2}$ QSLs using Schwinger boson mean-field theory, restricting our attention to QSLs which do not break any space group symmetries or time-reversal symmetry. We use the projective symmetry group (PSG) to classify all such QSLs [Wen2002, Wang2006, Liu2019]. We construct the phase diagram and identify the spin liquid candidates that are consistent with the experiments. Among them, the most robust candidates occur in the region where the transverse quadrupolar exchange coupling $J_{\pm}$ is dominant. This is a parameter regime where studies on the XXZ model (1) have identified a $U(1)$ nematic QSL through exact diagonalization and variational calculations [benton2018quantum] and a disordered state with a large nematic susceptibility in pseudo-Majorana fermion functional renormalization group [Schaden2025]. The $\mathbb{Z}_{2}$ QSLs identified here might be closely competing with these disordered states and may be stabilized by couplings beyond the pristine XXZ model. Finally, we compute the static and dynamic spin structure factors for these candidates to identify their experimental signatures in neutron scattering.
@@
Figure 1: (a) The pyrochlore lattice and global Cartesian coordinate frame, where “up-pointing” and “down-pointing” tetrahedra are coloured green and blue, respectively. (b) An up-pointing tetrahedron with local $\hat{z}$ axes and sublattice indices attached. (c) Phase diagram in the region $J_{\pm}<0$ with $J_{zz}^{2}+J_{\pm}^{2}+J_{\pm\pm}^{2}=1$ indicating the lowest energy PSG classes. We set $\kappa=0.1$. The XXZ model occurs along the horizontal line joining $\phi=0$ to $\phi=\pi$. The point H indicates the antiferromagnetic Heisenberg point $J_{\pm}/J_{zz}=-\frac{1}{2},J_{\pm\pm}=0$. The $\times$’s indicate the points at which the structure factors were calculated for the relevant PSG classes (see table 1). (d) Phase diagram at $\kappa=0.1$ indicating the second-lowest energy PSG classes.
@@ II Model
In praseodymium-based rare-earth pyrochlores, magnetic $\textrm{Pr}^{3+}$ ions are located at the shared vertices of neighbouring tetrahedra [Gardner2010]. Each $\textrm{Pr}^{3+}$ has two $f$-electrons, and an application of Hund’s rules results in a degenerate $J=4$ total angular momentum state. A crystal electric field (CEF) with $D_{3d}$ symmetry splits the $J=4$ manifold into doublets and singlets. Since there are an even number of electrons, Kramers’ theorem does not apply, so any degeneracy arises from crystal symmetries [Rau2019]. Experimentally, it is found that the lowest-lying states form a doublet (i.e., a non-Kramers doublet) and are well separated from the higher CEF states by at least $9$ meV [Machida2005, Martin2017, Sibille2018]. Thus, we are justified in keeping only the non-Kramers doublet degrees of freedom for a low-energy description of the physics. We denote the two degenerate states by $|\pm\rangle$ and define pseudospin-$1/2$ operators $S$ by
@@
$|\pm\rangle$ $=\alpha|m=\pm 4\rangle\pm\beta|m=\pm 1\rangle-\gamma|m=\mp 2\rangle$ (4)
where $\alpha,\beta,\gamma\in\mathbb{R}$ [Schaffer2013]. The resulting $S^{z}$ component of the non-Kramers doublet transforms as a dipole, while the in-plane $S^{x},S^{y}$ components transform as quadrupoles [Onoda2011].
@@
  • $S$: a “screw” operation defined by fractional translation by $\hat{\epsilon}_{3}=\frac{1}{2}{\bm{\hat{e}}}_{3}$ composed with a $\pi$ rotation around ${\bm{\hat{e}}}_{3}$. [Curnoe2018, Liu2019, Desrochers2022]
Finally, it is convenient to work in sublattice-dependent local bases to express the orientations of the pseudospins. We attach an orthonormal basis $\{\hat{x}_{\mu},\hat{y}_{\mu},\hat{z}_{\mu}\}$ to each sublattice $\mu$ (expressions given in appendix A). As shown in Fig. 1 (b), these have the property that each $\hat{z}_{\mu}$ basis vector points out of the “up-pointing” tetrahedra.
@@
where $i,j$ denote pyrochlore lattice sites, and $\gamma_{ij}$ is a bond-dependent phase factor defined in appendix A [Onoda2011]. We take $J_{\pm\pm}>0$ without loss of generality because its sign can be flipped by a $\frac{\pi}{2}$ rotation of each local basis $\{\hat{x}_{\mu},\hat{y}_{\mu},\hat{z}_{\mu}\}$ about $\hat{z}_{\mu}$ [Chung2024].
@@ III.1 Schwinger Bosons
$S^{a}_{i}$ $=\frac{1}{2}\sum_{\alpha\beta}b^{\dagger}_{i\alpha}[\sigma^{a}]_{\alpha\beta}b_{i\beta}$ (9)
where $a\in\{x,y,z\}$ and $\sigma^{a}$ are the Pauli matrices. The operators $b_{i\alpha}$ satisfy the bosonic commutation relations $b_{i\alpha},b^{\dagger}_{j\beta}]=\delta_{ij}\delta_{\alpha\beta}$. The Schwinger bosons physically represent the fractionalized excitations of the QSL, which we refer to as spinons. If we impose the local constraint
@@
with $\kappa=1$, we obtain a faithful representation of the original pseudospin-$1/2$ Hilbert space [Book2011, Wang2006]. In the mean-field theory, this constraint is relaxed to only be satisfied on average
@@
for all sites $i$. We will also treat $\kappa$ as a continuous positive parameter, which gives us control over the size of quantum effects. Taking $\kappa\to\infty$ recovers the classical limit, and $\kappa<1$ corresponds to larger quantum fluctuations. In the current work, we focus on small $\kappa\ll 1$ to obtain all the competing QSLs, and comment on what is expected in the physical $\kappa=1$ limit later in sections IV.2 and V [Messio2010]. As $\kappa$ increases, magnetic order may develop in some parts of the phase diagram, signaled by the closing of the spinon gap and condensation of Schwinger bosons [Wang2006, Dey2024, Liu2019].
@@ III.2 Mean-Field Decoupling
$H_{\textrm{MF}}$ $=-\sum_{\langle ij\rangle}\left(\bm{b}_{i}^{\dagger}\left[u^{h}_{ij}\right]\bm{b}_{j}+\bm{b}_{i}^{\intercal}\left[u^{p}_{ij}\right]\bm{b}_{j}\right)+\textrm{h.c.}$
@@ III.3 Projective Symmetry Group Classification
Two ansätze which differ by a gauge transformation of the form (19) are therefore physically equivalent [Wen2002, wen2004quantum, Wang2006]. The correct way to implement symmetry operations is via the gauge-enriched operations $\tilde{{\mathcal{O}}}=G_{\mathcal{O}}\circ{\mathcal{O}}$, where $\mathcal{O}$} is an element of the pyrochlore lattice space group [Wen2002, Wang2006]. Explicitly, the action of $\tilde{{\mathcal{O}}}$ is
@@
with $\mathcal{K}$ being the complex conjugation operator and $U_{\mathcal{T}}$ is a $2\times 2$ unitary matrix [Liu2019]. Explicit expressions for $U_{\mathcal{O}}$ and $U_{\mathcal{T}}$ are derived in appendix B; here we simply list the results
@@
where $\omega=e^{2\pi i/3}$. The group generated by the gauge-enriched operations $\tilde{{\mathcal{O}}}$ and $\tilde{\mathcal{T}}$ is known as the projective symmetry group (PSG) [Wen2002, Wang2006, Liu2019]. Demanding that the mean-field Hamiltonian is invariant under the PSG elements puts constraints on the allowed mean-field ansätze. Ansätze that cannot be transformed into each other via a pure gauge transformation (19) belong to different PSG classes, meaning they describe different QSLs [Wen2002, Wang2006].
The subgroup of the PSG corresponding to pure gauge transformations associated with the identity operation ${\mathcal{O}}=\mathrm{id}$ is known as the invariant gauge group [Wen2002, Wang2006]. The invariant gauge group corresponds to the emergent gauge structure of the resulting QSLs [Wen2002, wen2004quantum]. In this paper, we only consider the $\mathbb{Z}_{2}=\{e^{i\pi n}|n\in\{0,1\}\}$ invariant gauge group, thus restricting our study to $\mathbb{Z}_{2}$ symmetric bosonic QSLs.
@@
where $n_{1},n_{\overline{C}_{6}S},n_{ST_{1}},n_{\overline{C}_{6}}$ are all either 0 or $1$ [Liu2019]. Even though the PSG solution is the same as the effective spin-$1/2$ [Liu2019] and dipolar-octupolar cases [Desrochers2022], the structure of the allowed mean-field ansätze will be different because of the pseudospin transformation matrices $U_{\mathcal{O}}$ and $U_{\mathcal{T}}$. Following [Liu2019], we shift the first Brillouin zone ${{\bm{q}}}\mapsto{{\bm{q}}}-(\pi,\pi,\pi)$ for the PSG classes with ${n_{ST_{1}}}=1$, which is equivalent to a gauge transformation and does not affect physical quantities on the spin level.
The PSG classes with $n_{1}=1$ have an enlarged unit cell because the translations $T_{2}$ and $T_{3}$ act projectively, which means that the translation-invariance of the ansatz is realized after projecting from the Schwinger boson Hilbert space down to the physical spin Hilbert space. We do not consider these classes in this work. There is no physical reason to exclude them, but only considering $n_{1}=0$ simplifies the calculations and the presentation of results. We therefore label the different $n_{1}=0$ PSG classes using the notation $(n_{\overline{C}_{6}S}n_{ST_{1}}n_{\overline{C}_{6}})$.